OpenStudy (anonymous):

Solve by completing the square. A. {2, 6} B. {2, –6} C. {–2, 6} D. {–2, –6} Please select the best answer from the choices provided

OpenStudy (anonymous):

S^2-8s+12=0

OpenStudy (amistre64):

what would complete the square on the "s" parts?

OpenStudy (amistre64):

s^2 -8s + _____

OpenStudy (anonymous):

12

OpenStudy (amistre64):

no

OpenStudy (amistre64):

the 12 is extra baggage at the moment

OpenStudy (anonymous):

Would I have to factor the -8?

OpenStudy (amistre64):

a complete square is of the form: (s+n)^2 = s^2 + 2ns + n^2 if we compare coeffs s^2 + 2ns + n^2 = s^2 - 8s + _____ , then what does "n" have to be? and thereby n^2?

OpenStudy (anonymous):

-4?

OpenStudy (amistre64):

correct, and (-4)^2 = 16 lets add zero to your set up; since 16-16 = 0 ... s^2 -8s +16 - 16 +12 = 0

OpenStudy (amistre64):

now we can "compact" the square into its (s+n)^2 format (s-4)^2 -4 = 0

OpenStudy (anonymous):

Then add -4 to both sides?

OpenStudy (amistre64):

umm, lets just add +4 to each side :) but yes

OpenStudy (anonymous):

That's what I meant lol (s-4)^2=4

OpenStudy (amistre64):

good, now lets sqrt, and solve for s

OpenStudy (anonymous):

The square root would be 2?

OpenStudy (amistre64):

there are two values that when squared equal 4; (2)^2 = 4 and (-2)^2 = 4 always remember to include a +- when sqrting

OpenStudy (anonymous):

So (x+2)=4 (x+2)=-4. then subtract 2 from both sides?

OpenStudy (anonymous):

x=2 and x =-6?

OpenStudy (amistre64):

let me clean that up aliitle \[(s-4)^2=4\] \[\sqrt{(s-4)^2}=\pm\sqrt{4}\] \[s-4=\pm2\] \[s-4+4=4\pm2\] \[s=4\pm2~;~2,6\]

OpenStudy (anonymous):

Thank you, it helped a lot. I'd rather have help on it than just get the answer.

OpenStudy (amistre64):

youre welcome, and good luck :)